OpenStudy (anonymous):

x^2+kx+k+1=0 a) for what values of k will the equation have two real roots? b) fin the largest whole number value of k for which the equation will have two complex solution c) solve the equation for the value of k found in part (b)

OpenStudy (aravindg):

easy

OpenStudy (aravindg):

discriminant should be greater than o0 to have 2 roots ..equate this and u get value of k

OpenStudy (aravindg):

i see k can can have four values here

OpenStudy (aravindg):

see here

OpenStudy (aravindg):

all efforts and no medal?

OpenStudy (anonymous):

\[\Delta = k^2 -4(1)(k+1)\]

OpenStudy (anonymous):

Now we need this discriminant to be POSITIVE for two real roots. It is a quadratic expression so check the discriminant of the discriminant that we just calculated. \[\Delta2= (-4)^2 -4(1)(-4) = 16+16 >0 \]

OpenStudy (anonymous):

So the original eqn has two real roots for all k. That can't seem correct given it has extra parts.

OpenStudy (anonymous):

In this case a=1,b=k,c=k+1, \[\Delta=k ^{2}-4\left( k +1 \right)=k^{2}-4k -4\]\[k=\left( 4\pm \sqrt{16-4\times -4} \right)\div2\]\[k=\left( 4\pm \sqrt{32} \right)\div2=\left( 4\pm4\sqrt{2} \right)\div2=2\pm2\]